OpenStudy (anonymous):

Suppose that to make the golf team you need to score no more than 74 on average over 5 games. If you scored 77, 71, 77, and 67 in your first 4 games what is the highest score you can shoot in your 5th game and still make the team? A. 76 B. 78 C. 79 D. 80

OpenStudy (anonymous):

D?

OpenStudy (calculusxy):

\[\large \frac{ t_1 + t_2 + t_3 + t_4 + t_5 }{ 5 } = \text{average}\]

OpenStudy (anonymous):

how did u get the answer

OpenStudy (calculusxy):

\(\large t_1\) to \(\large t_4\) are values that you can replace with 77, 71, 77, and 67. Since we are trying to find \(t_5\), we can leave it as it is. I put the equation over 5 because we will be taking the average of 5 games. And for the part that says "average" we can replace it with 74.

OpenStudy (calculusxy):

\[\large \frac{ 77 + 71 + 77 + 67 + t_5 }{ 5 } = 74\] Do you know how to solve this?

OpenStudy (calculusxy):

Meaning solve for \(\large t_5\)?

OpenStudy (calculusxy):

@vduverge26 Did you find the answer or do you need help?

OpenStudy (anonymous):

yes thank u